solve for x , 2x(x-3)(x-4)=0
Answer by malakumar_kos@yahoo.com(315)


2x(x-3)(x-4)=0 solve for x
From the given expression we get 2x = 0 or x = 0
x-3 = 0 or x = 3
x-4 = 0 or x = 4
The solution for x = 0,3,4